# Adaptive Exploration for Multi-Reward Multi-Policy Evaluation

### v2

@@ B.2.2 Optimality of Behavior Policies
$\sup_{r\in{\cal R}}\max_{s}\|\rho_{r}^{\pi}(s)\|_{\infty}^{2}$ $=\max_{s}\sup_{r\in{\cal R}}\max_{s^{\prime}}|\rho_{r}^{\pi}(s,s^{\prime})|^{2},$
$=\max_{s}\max_{s^{\prime}}\left(\sup_{r\in{\cal R}}|\rho_{r}^{\pi}(s,s^{\prime})|\right)^{2},$
$=\max_{s}\max_{s^{\prime}}\max\left(\sup_{r\in{\cal R}}\rho_{r}^{\pi}(s,s^{\prime}),\sup_{r\in{\cal R}}-\rho_{r}^{\pi}(s,s^{\prime})\right)^{2}.$
@@
In the last part of the proof we consider the case where ${\cal R}=[0,1]^{S}$. We exploit the following fact that holds for any vector $y\in\mathbb{R}^{n}$:$\max_{x\in[0,1]^{n}}y^{\top}x=\sum_{i:y_{i}>0}y_{i}$. Then, let $\Gamma_{ij}^{\pi}(s)\coloneqq\left(K^{\pi}(s)G^{\pi}\right)_{ij}$ and define
@@ B.2.5 Computing the Optimal Relaxed Characteristic Rate $U^{\star}$
Therefore, given a collection of deterministic target policies $\{\pi_{i}\}_{i=1}^{N}$ and reward sets $\{\mathcal{R}_{i}\}_{i=1}^{N}$ (finite, or a convex set), we pre–compute, for every policy $i$ and state $s$,
@@
The inner sup is solved twice in the general case. For a finite reward set it is a direct maximum, and one can directly compute $A_{i}(s)=\max_{s^{\prime}}\max_{r\in{\cal R}_{i}}|\rho_{r}^{\pi}(s,s^{\prime})|^{2}$. For the reward free case we can use the result from Corollary 4.7 to obtain $A_{i}(s)=\max_{s^{\prime}}\max\left(\Gamma_{+}^{\pi}(s,s^{\prime}),\Gamma_{-}^{\pi}(s,s^{\prime})\right)^{2}$.
Stacking the rows $A_{i}$ yields a matrix $A\in\mathbb{R}^{N\times|\mathcal{S}|}$. For a stationary occupancy measure $\omega\in\Omega(M)$ we define $f_{i}(\omega)=\displaystyle\max_{s\in\mathcal{S}}A_{i}(s)\,\frac{\gamma^{2}}{4\epsilon^{2}(1-\gamma)^{2}}\,\frac{1}{\omega(s,\pi_{i}(s))}.$ The optimal relaxed rate is obtained from the convex optimisation problem
@@
Figure 5: Complexity of Riverswim for varying reward sets and state space sizes for a single policy $\pi(\cdot)=a_{2}$ that moves towards the river’s end. Left: complexity for different $p$ values (prob. of moving toward the river’s end). Right: complexity for $p=0.3$ with varying state space sizes. Solid curves evaluate on ${\cal R}=[0,1]^{S}$, dashed curves on =${\cal R}=\{r\in[0,1]^{S}:\|r\|_{2}\leq 1\}$.
@@ Appendix C MR-NaS and Numerical Results
Let ${\cal C}_{\pi,r}=\{\|\hat{V}_{r}^{\pi}-V_{r}^{\pi}\|_{\infty}\geq\epsilon\}$ be the event that at the stopping time the value estimate is not $\epsilon$-accurate in $\pi,r\in{\cal R}_{\pi}$. Hence, we have that $M\in{\rm Alt}_{\pi,r}^{\epsilon/2}(M_{\tau})$.
@@
${\cal C}=\{\exists\pi\in\Pi,r\in{\cal R}_{\pi}:\|\hat{V}_{r}^{\pi}-V_{r}^{\pi}\|_{\infty}\geq\epsilon\}\subset\{\exists\pi\in\Pi,r\in{\cal R}_{\pi}:M\in{\rm Alt}_{\pi,r}^{\epsilon/2}(M_{\tau})\}\coloneqq{\cal B}.$
@@
The sample complexity results follow from noting that $U_{\epsilon/2}(\omega;M)=4U_{\epsilon}(\omega;M)$ and applying the same methods as in Russo & Vannella (2024, Theorem 3.3) mutatis mutandis. ∎
@@ C.2 Environment Details
In figure Figure 6 is shown a depiction of the environment. There are $n$ states, and two main parameters, $p,p^{\prime}\in(0,1)$, and their sum $p_{\rm sum}=p+p^{\prime}<1$. In the figure, each tuple $(a,p)$ represents the action $a$ that triggers the transition and the probability $p$ of that event. The agent starts in state $s_{0}$, and in every state can only take two actions $\{a_{0},a_{1}\}$. For small values of $p$ it becomes difficult for the agent to swim right (i.e., reach $s_{n-1}$), and larger values of $p^{\prime}$ can also hinder the progress. On the other hand, swimming towards the left is easier, since the probability of $P(s_{i-1}|s_{i},a_{0})=1$. For the experiments, we used $n\in\{15,20,30\},p=0.7,p^{\prime}=6(1-p)/7$
@@
In Figure 7 is shown a depiction of the environment. There are a total of $12n+$ 1 states, and two parameters $p,p^{\prime}\in(0,1)$, so that $p_{\rm sum}=p+p^{\prime}<1$. In the figure, each tuple $(a,p)$ represents the action $a$ that triggers the transition and the probability $p$ of that event. The agent starts in state $s_{0}$, and in every state can chose between three actions $\{a_{0},a_{1},a_{2}\}$. For small values of $p$ it becomes difficult for the agent to swim right in both forks, and larger values of $p^{\prime}$ can also hinder the progress. As in Riverswim, swimming towards the left is easier, since the probability of $P(s_{i-1}|s_{i},a_{0})=1$. For the experiments, we used $n\in\{8,10,15\},p=0.7,p^{\prime}=6(1-p)/7$.
@@
In Figure 8 is shown a depiction of the environment. There are a total of $12n+$ 1 states, and one parameters $p\in(0,1)$. In the figure, each tuple $(a,p)$ represents the action $a$ that triggers the transition and the probability $p$ of that event. The agent starts in state $s_{0}$, and in every state can chose between two actions $\{a_{0},a_{1}\}$. For small values of $p$ it becomes difficult for the agent to move to the end of the chain in both chains. For the experiments, we used $n\in\{8,10,15\},p=0.7$.
@@ C.4 Additional Results and Experimental Details
In the finite reward‐set scenario, we restrict =${\cal R}=\{r_{i}\}_{i}$ to a uniformly chosen subset of size $3$ from the canonical basis ${\cal R}_{\rm canon}$ (sampled without replacement).

### v3

@@ B.2.2 Optimality of Behavior Policies
$\sup_{r\in{\cal R}_{\pi}}\max_{s}\|\rho_{r}^{\pi}(s)\|_{\infty}^{2}$ $=\max_{s}\sup_{r\in{\cal R}_{\pi}}\max_{s^{\prime}}|\rho_{r}^{\pi}(s,s^{\prime})|^{2},$
$=\max_{s}\max_{s^{\prime}}\left(\sup_{r\in{\cal R}_{\pi}}|\rho_{r}^{\pi}(s,s^{\prime})|\right)^{2},$
$=\max_{s}\max_{s^{\prime}}\max\left(\sup_{r\in{\cal R}_{\pi}}\rho_{r}^{\pi}(s,s^{\prime}),\sup_{r\in{\cal R}_{\pi}}-\rho_{r}^{\pi}(s,s^{\prime})\right)^{2}.$
@@
In the last part of the proof we consider the case where ${\cal R}_{\pi}=[0,1]^{S}$. We exploit the following fact that holds for any vector $y\in\mathbb{R}^{n}$$\max_{x\in[0,1]^{n}}y^{\top}x=\sum_{i:y_{i}>0}y_{i}$. Then, let $\Gamma_{ij}^{\pi}(s)\coloneqq\left(K^{\pi}(s)G^{\pi}\right)_{ij}$ and define
@@ B.2.5 Computing the Optimal Relaxed Characteristic Rate $U^{\star}$
Therefore, given a collection of deterministic target policies $\{\pi_{i}\}_{i=1}^{N}$ and reward sets $\{\mathcal{R}_{i}\}_{i=1}^{N}$ (finite, or a convex set), we precompute, for every policy $i$ and state $s$,
@@
The inner sup is solved twice in the general case. For a finite reward set it is a direct maximum, and one can directly compute $A_{i}(s)=\max_{s^{\prime}}\max_{r\in{\cal R}_{i}}|\rho_{r}^{\pi}(s,s^{\prime})|^{2}$. For the reward free case we can use the result from Corollary 4.8 to obtain $A_{i}(s)=\max_{s^{\prime}}\max\left(\Gamma_{+}^{\pi_{i}}(s,s^{\prime}),\Gamma_{-}^{\pi_{i}}(s,s^{\prime})\right)^{2}$.
Stacking the rows $A_{i}$ yields a matrix $A\in\mathbb{R}^{N\times|\mathcal{S}|}$. For a stationary occupancy measure $\omega\in\Omega(M)$ we define $f_{i}(\omega)=\displaystyle\max_{s\in\mathcal{S}}A_{i}(s)\,\frac{\gamma^{2}}{2\epsilon^{2}(1-\gamma)^{2}}\,\frac{1}{\omega(s,\pi_{i}(s))}.$ The optimal relaxed rate is obtained from the convex optimisation problem
@@
Figure 5: Complexity of Riverswim for varying reward sets and state space sizes for a single policy $\pi(\cdot)=a_{2}$ that moves towards the river’s end. Left: complexity for different $p$ values (prob. of moving toward the river’s end). Right: complexity for $p=0.3$ with varying state space sizes. Solid curves evaluate on ${\cal R}=[0,1]^{S}$, dashed curves on ${\cal R}=\{r\in[0,1]^{S}:\|r\|_{2}\leq 1\}$.
@@ Appendix C MR-NaS and Numerical Results
Let ${\cal C}_{\pi,r}=\{\|\hat{V}_{r}^{\pi}-V_{r}^{\pi}\|_{\infty}>\epsilon\}$ be the event that at the stopping time the value estimate is not $\epsilon$-accurate in $\pi,r\in{\cal R}_{\pi}$. Hence, we have that $M\in{\rm Alt}_{\pi,r}^{\epsilon/2}(M_{\tau})$.
@@
${\cal C}=\{\exists\pi\in\Pi,r\in{\cal R}_{\pi}:\|\hat{V}_{r}^{\pi}-V_{r}^{\pi}\|_{\infty}>\epsilon\}\subset\{\exists\pi\in\Pi,r\in{\cal R}_{\pi}:M\in{\rm Alt}_{\pi,r}^{\epsilon/2}(M_{\tau})\}\coloneqq{\cal B}.$
@@
For the sample complexity, we use the following facts:
  1. 1.
    The forcing policy is chosen as $\pi_{f,t}(\cdot|s)=\texttt{softmax}\left(-\beta_{t}(s)N_{t}(s,\cdot)\right)$ with $\beta_{t}(s)=\frac{\beta\log(N_{t}(s))}{\max(1,\max_{a}|N_{t}(s,a)-\min_{b}N_{t}(s,b)|)},\beta\in[0,1]$ and $({\tt softmax}(x))_{i}=e^{x_{i}}/\sum_{j}e^{x_{j}}$ for a vector $x$. This choice encourages to select under-sampled actions for $\beta>0$, while for $\beta=0$ we obtain a uniform forcing policy $\pi_{f,t}(a|s)=1/A$. We then mix $\omega_{t}^{\star}$ with $\pi_{f,t}$ using a mixing factor $\epsilon_{t}=1/\max(1,N_{t}(s_{t}))^{\alpha}$, with $N_{t}(s)=\sum_{a}N_{t}(s,a)$. The values $\alpha,\beta$ need to guarantee $\alpha+\beta\leq 1$ (Russo & Vannella, 2024). Under this choice the chain induced by $\pi_{t}$ is ergodic.
  2. 2.
    We also note that the optimal solution $\omega^{\star}$ satisfies $\omega^{\star}(s,a)>0$ for all $(s,a)$. Then, such solution induces an ergodic (irreducible and aperiodic) chain by 3.1 and 5.1.
  3. 3.
    We note that the solution $\omega^{\star}$ is unique by 5.1.
  4. 4.
    Then, the sample complexity results follow from noting that $U_{\epsilon/2}(\omega;M)=4U_{\epsilon}(\omega;M)$ and applying the same methods as in Russo & Vannella (2024, Theorem 3.3) mutatis mutandis (which follows the proof technique of (Al Marjani et al., 2021)).
@@ C.2 Environment Details
In figure Figure 6 is shown a depiction of the environment. There are $n$ states, and two main parameters, $p,p^{\prime}\in(0,1)$, and their sum $p_{\rm sum}=p+p^{\prime}<1$. In the figure, each tuple $(a,p)$ represents the action $a$ that triggers the transition and the probability $p$ of that event. The agent starts in state $s_{0}$, and in every state can only take two actions $\{a_{0},a_{1}\}$. For small values of $p$ it becomes difficult for the agent to swim right (i.e., reach $s_{n-1}$), and larger values of $p^{\prime}$ can also hinder the progress. On the other hand, swimming towards the left is easier, since the probability of $P(s_{i-1}|s_{i},a_{0})=1$. For the experiments, we used $n\in\{15,20,30\},p=0.7,p^{\prime}=6(1-p)/7$ .
@@
In Figure 7 is shown a depiction of the environment. There are a total of $2n+1$ states, and two parameters $p,p^{\prime}\in(0,1)$, so that $p_{\rm sum}=p+p^{\prime}<1$. In the figure, each tuple $(a,p)$ represents the action $a$ that triggers the transition and the probability $p$ of that event. The agent starts in state $s_{0}$, and in every state can chose between three actions $\{a_{0},a_{1},a_{2}\}$. For small values of $p$ it becomes difficult for the agent to swim right in both forks, and larger values of $p^{\prime}$ can also hinder the progress. As in Riverswim, swimming towards the left is easier, since the probability of $P(s_{i-1}|s_{i},a_{0})=1$. For the experiments, we used $n\in\{8,10,15\},p=0.7,p^{\prime}=6(1-p)/7$.
@@
In Figure 8 is shown a depiction of the environment. There are a total of $2n+1$ states, and one parameters $p\in(0,1)$. In the figure, each tuple $(a,p)$ represents the action $a$ that triggers the transition and the probability $p$ of that event. The agent starts in state $s_{0}$, and in every state can chose between two actions $\{a_{0},a_{1}\}$. For small values of $p$ it becomes difficult for the agent to move to the end of the chain in both chains. For the experiments, we used $n\in\{8,10,15\},p=0.7$.
@@ C.4 Additional Results and Experimental Details
In the finite reward‐set scenario, we restrict ${\cal R}=\{r_{i}\}_{i}$ to a uniformly chosen subset of size $3$ from the canonical basis ${\cal R}_{\rm canon}$ (sampled without replacement).